OpenStudy (anonymous):

Use trigonometric identities to transform the left side of the equation into the right side (0 < θ < π/2). sin2 θ − cos2 θ=sin2 θ − (?????) = 2 sin2 θ − 1

OpenStudy (campbell_st):

well \[\sin^2 + \cos^2 = 1\] rewrite this equation \[\cos^2 (x) = 1- \sin^2(x)\] substitute into the original equation \[\sin^2(x) - (1 - \sin^2(x) = 2\sin^2(x) - 1\]

OpenStudy (anonymous):

i just need to know what to put in the ???? part. this is confusing me. i understand what you mean though.

OpenStudy (anonymous):

Just looking at it (no calculation, formula or anything) if you put (1-sin2 theta) then you get the last line.

OpenStudy (campbell_st):

then you need to solve 2sin^2(x) = 1 sin^2x = 1/2 and \[\sin(x) = \pm \frac {1}{\sqrt{2}}\]

OpenStudy (campbell_st):

1st quadrant as 0<x<pi/2 then x = pi/4

OpenStudy (anonymous):

The questioner is not actually asking for anything to be solved.....

OpenStudy (campbell_st):

I substituted 1 - sin^2(x) for cos^2(x)